OpenStudy (anonymous):

Given the functions f(x) = 7x + 13 and g(x) = x + 2, which of the following functions represents f[g(x)] correctly?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

f[g(x)] = 7x + 15 f[g(x)] = 7x + 27 f[g(x)] = 7x2 + 15 f[g(x)] = 7x2 + 27

jimthompson5910 (jim_thompson5910):

f(x) = 7x+13 f(x) = 7( x )+13 f(g(x)) = 7( g(x) )+13 ... replace every x with g(x) f(g(x)) = 7( x+2 )+13 ... replace every g(x) on the right side with x+2

jimthompson5910 (jim_thompson5910):

I'll let you simplify

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

7x+27?

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

ok thank u very much

jimthompson5910 (jim_thompson5910):

no problem